# Maintaining Distributed Data Structures in Dynamic Peer-to-Peer Networks

### v1

@@ 2.4 Buffer Creation
In this section, we outline the process of constructing the Buffer network $\mathcal{B}$, which comprises the nodes that have joined the network. Conceptually, to establish a skip list from an unsorted collection of distinct elements, the initial step involves creating the lowest level by organizing a sorted list. Subsequently, for each node in this list, a random experiment is conducted to construct the upper levels. In our context, we encounter an unsorted set of nodes $C$, which have been introduced into the Spartan network $\mathcal{S}$ through adversarial churn, and our objective is to construct a distributed skip list from it. Initially, our challenge lies in efficiently arranging these nodes to form the lowest level of the distributed data structure. The current best-known approach to sort $n$ elements using $\text{polylog}{(n)}$ incoming and outgoing messages for each node during each round requires $\mathcal{O}(\log^{3}n)$ time [6]. To improve on such result, we rely on sorting networks theory (e.g. [16, 1, 53]). A sorting network is a graph topology which is carefully manufactured for sorting. More precisely, it can be viewed as a circuit-looking DAG (directed acyclic graph) with $n$ inputs and $n$ outputs in which each node has exactly two inputs and two outputs (i.e., two incoming edges and two outgoing edges). A sorting network is such that for any input, the output is monotonically sorted. A sorting network is characterized by two parameters: the number of nodes resp. size and the depth. The space required to store a sorting network is determined by its size while the overall running time of the sorting algorithm is determined by its depth. Figure 6 shows an example of sorting network on $4$ elements.
@@
Figure 6: Example of a sorting network [16] on $4$ numbers. Every crossing edge connecting two nodes in different levels is a node in the DAG representation. The upper output of a crossing edge connecting two nodes in different levels is the minimum between the two inputs, while the lower output is the maximum. Execution of the sorting network is left to right.
One notable construction for sorting $n$ elements is the $\mathcal{O}(\log n)$-depth (thus $\mathcal{O}(n\log n)$ size) AKS network, developed by Ajtai, Komlós, and Szemerédi [1] in the early $80$’s. However, a drawback of the AKS network is that the $\mathcal{O}(\cdot)$ notation hides large constant factors. Some progress has been made in simplifying the AKS network and improving the constant factors in its depth [66], but for practical values of $n$, the depth of Batcher’s bitonic sort [16] remains considerably smaller. Nevertheless, if we disregard the impracticability of the AKS network, we can improve the distributed sorting result in [6] by a $\mathcal{O}(\log^{2}n)$ factor. Moreover, it is well known that sorting networks can be simulated by a butterfly network and other hypercubic networks (see Chapter 3.5 in [53]). So, we first need to find a way to embed the AKS sorting network on a butterfly-like one. To this end, we use the result by Maggs and Vöcking [56] in which they show that an AKS network can be embedded on a multibutterfly [84]. A $k$-dimensional multibutterfly network consists of $k+1$ levels, each consisting of $2^{k}$ nodes. For $0\leq\ell\leq k$ and $0\leq i<2^{k}$, let $(\ell,i)$ be the label of the $\ell$th node on level $i$. The nodes on each level $\ell$ are partitioned into $2^{k}$ sets $A_{\ell,0},\dots,A_{\ell,2^{k}-1}$ where $A_{\ell,j}=\{(\ell,i):\lfloor i/2^{k-\ell}\rfloor=j\}$. The nodes in $A_{\ell,j}$ are connected to the nodes in $A_{\ell+1,2j}$ and $A_{\ell+1,2j+1}$. To embed the AKS network we need to consider a subclass of the multibutterfly networks that includes those multibutterfiles that can be constructed by superimposing butterfly networks. Furthermore, Maggs and Vöcking [56] showed how to embed the AKS network on what they called twinbutterfly, a multibutterfly obtained by superimposing two butterfly networks. On the twinbutterflty each leaf node has degree $4$ and each non-leaf node has degree $8$.
For completeness we state the result we are interested in.
@@
Here the load of an embedding is the maximum number of nodes of the AKS network mapped to any node of the twinbutterfly. The congestion is the maximum number of paths that use any edge in the twinbutterfly, and the dilation is the length of the longest path in the embedding. It follows that
@@
The idea is to use such a butterfly construction algorithm to build a twinbutterfly $\mathcal{M}$ in $\mathcal{O}(\log n)$ rounds and simulate the AKS algorithm on it. Moreover, $\mathcal{M}$ being a data structure build on top of the Spartan network, it naturally enjoys a churn resiliency property. In fact, every time a node $v$ is churned out from $\mathcal{S}$, a committee will temporary cover for $v$ in $\mathcal{M}$ (see Section 2.2). Finally, we show that the $\mathcal{B}$ can be created in $\mathcal{O}(\log n)$ rounds.
@@
$\mathcal{M}$ can be built in $\mathcal{O}(\log n)$ rounds (Lemma 2.6). Moreover, to handle cases in which $\kappa\cdot N+o(N)$ (see Theorem 2.5) is greater than the number of nodes in the Spartan network $\mathcal{S}$, we allow for nodes to represent a constant number of temporary “dummy” nodes for the sake of building the needed multibutterfly. Once $\mathcal{M}$ has been constructed, we run AKS on such network and after $\mathcal{O}(\log n_{\mathcal{B}})=\mathcal{O}(\log n)$ rounds we obtain the sorted list at the base of the new skip list we want to build. Next, each node $u$ in the buffer computes its maximum height $\ell_{\max}^{\mathcal{B}}(u)$ in the buffer skip list $\mathcal{B}$ and each level $0<\ell\leq\ell_{\max}(\mathcal{B})$ is created by copying the base level of the skip list. Moreover, each node $u$ at level $0<\ell\leq\ell_{\max}(\mathcal{B})$ can be of two types, effective or fill-in. We say that a node $u$ is effective at level $\ell$ if $\ell\leq\ell_{\max}^{\mathcal{B}}(u)$, and fill-in otherwise (see Figure 7(a)). Next, we use the same parallel $\mathcal{O}(\log n)$ rounds rewiring technique used in the Deletion phase (see Section 2.3) to exclude fill-in nodes from each level and obtain a skip list of effective nodes (see Figures 7(b)-7(c)). Wrapping up, to create the buffer network we need $\mathcal{O}(\log n)$ rounds to build the twibutterfly $\mathcal{M}$, $\mathcal{O}(\log n)$ rounds to run AKS on $\mathcal{M}$, $\mathcal{O}(\log n)$ rounds w.h.p. to duplicate the levels (see Lemma B.1 in Appendix B) and $\mathcal{O}(\log n)$ rounds to run the rewiring procedure. Thus, we need $\mathcal{O}(\log n)$ rounds [^8] We point out that the constant hidden by the $\mathcal{O}(\cdot)$ notation is the one for the AKS sorting network [66]. w.h.p.
@@ 2.5 Merge
As a starting point, we analyze the process of inserting (in parallel) a sequence of sorted elements $X$ in a skip list $\mathcal{C}$. More precisely, each element $v\in X$ starts the insertion process in $\mathcal{C}$ at the same time and from the same spot. Intuitively, all these parallel insertions can be seen as a group of elements traveling all together in the skip list. We will use the term cohesive group to describe a group of elements traversing the skip list together. We define the leader of the cohesive group to be the element with the smallest ID/key among the group. In addition, the group is assumed to be small enough to ensure all nodes can exchange messages pairwise simultaneously in one round. While traversing the skip list, the leader is in charge of querying the newly encountered nodes and to communicate their ID/key to the cohesive group. Furthermore, it can happen that at some point a subset $X^{\prime}\subseteq X$ separates from $X$ and takes another route to level 00. When such a split happens, say $X$ splits in two groups (see Proposition 2.10 below), the newly created cohesive group $X^{\prime}$ elects as a leader the element with the smallest ID/key in $X^{\prime}$. Note that such a split requires two rounds, one for the leader to inform all nodes in the new cohesive group who their leader is and one for the nodes in the new cohesive group to “introduce” themselves to their new leader.
Proposition 2.10.
Let $X=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{i-1}\leq x_{i}$ for $i\in[1,k]$ be a set of nodes that are traversing all together (i.e., a cohesive group) the skip list $\mathcal{C}$. And let $v\in X$ be a node that separates from $X$ at some time $t\geq 0$, then all $u\in X$ such that $u\geq v$ will separate from $X$ as well and will follow $v$.
@@
For the sake of contradiction, let $X=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{i-1}\leq x_{i}$ for $i\in[1,k]$ be a cohesive group that is traversing the skip list. Moreover, assume that $x_{j}$ separates from $X$ at some time $t$, and that there exists $x_{\ell}$ for $\ell\geq j$ that does not follow $x_{j}$, i.e., $x_{\ell}$ does not separate from $X$. This generates a contradiction on the assumption that the elements in $X$ are sorted in ascending order. Thus, it can not happen.
Finally, we show that the time to insert a cohesive group in a skip list is at most twice the time to insert a single element in the skip list.
Lemma 2.12.
Let $X=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{i-1}\leq x_{i}$ for $i\in[1,k]$ be a set of cohesive nodes that are inserted all together at the same time and from the top left sentinel node in a skip list $\mathcal{C}$. Then, the time to insert $x_{k}$ along with $x_{0},x_{1},\dots,x_{k-1},x_{k}$ is at most twice the time to insert just $x_{k}$.
@@
[…]
WAVE: An $\bm{\mathcal{O}(\log n)}$ merge procedure.
Here we show one of the main contributions of this paper. We provide a distributed protocol to merge two skip lists in $\mathcal{O}(\log n)$ rounds w.h.p. Intuitively, our merge procedure can be viewed as a top-down wave of parallel searches. The searches are performed on $\mathcal{C}$ by elements in $\mathcal{B}$. The wave is initiated by the top-most nodes in $\mathcal{B}$. At any point in time, the wave front represents all the nodes in $\mathcal{B}$ that are actively looking for the right position in $\mathcal{C}$. As nodes on the wave front find their respective appropriate location, they get inserted into $\mathcal{C}$. Thus, when the wave has swept through $\mathcal{C}$ in its entirety, we get the required merged skip list. Moreover, we notice that there must be only one cohesive group in the top level of the skip list $\mathcal{B}$, and that each couple of cohesive groups $X_{\ell},Y_{\ell}$ at the same level $\ell<\ell_{\max}(\mathcal{B})$ must be separated by one (or more) nodes $v$ that belongs to one (or more) cohesive group $Z_{\ell^{\prime}}$ such that $\ell^{\prime}>\ell$.
@@
Let $\ell$ be the maximum height of the skip list $\mathcal{B}$ and assume to have two disjoint cohesive groups $X_{\ell}$ and $Y_{\ell}$. This implies that $X_{\ell}$ and $Y_{\ell}$ are separated by one (or more nodes) $v$ such that $\ell_{\max}^{\mathcal{B}}(v)>\ell$. Thus we have a contradiction on $\ell$ being $\mathcal{B}$’s maximum height. Next, let $0\leq\ell<\ell_{\max}(\mathcal{B})$, consider the cohesive group $X_{\ell}=\{x_{0},x_{1},\dots,x_{k}\}$ and its left neighbor $v$ at level $\ell$ (i.e., $v\in N_{\mathcal{B}}^{(L,\ell)}(x_{0})$). Then $v$ must be such that $\ell_{\max}^{\mathcal{B}}(v)>\ell$. That is because, if $\ell_{\max}^{\mathcal{B}}(v)=\ell$ then we would have $v\in X_{\ell}$ and if $\ell_{\max}^{\mathcal{B}}(v)<\ell$, we would have $v\notin N_{\mathcal{B}}^{(L,\ell)}(x_{0})$.
Before diving into the description of the WAVE protocol, we define the left and right children of a node $u$ at some level $\ell>0$ in a skip list. Informally, the right (resp. left) children of a node $u$ at level $\ell$ is defined by the set of contiguous nodes after (resp. before) $u$ at level $\ell$ in the skip list. Formally, given a node $u$ in the buffer network $\mathcal{B}$ at level $\ell>0$ we define $\Gamma_{\ell^{\prime}}^{L}(u)=\{v:v<u\land\ell_{\max}(v)=\ell^{\prime}\land\forall v<w<u,\ell_{\max}(w)\leq\ell^{\prime}\}$and $\Gamma_{\ell^{\prime}}^{R}(u)=\{v:v>u\land\ell_{\max}(v)=\ell^{\prime}\land\forall u<w<v,\ell_{\max}(w)\leq\ell^{\prime}\}$ as the left and right children of $u$ at level $\ell^{\prime}<\ell$ in $\mathcal{B}$. Moreover, we refer to $u$’s set of children at level $\ell^{\prime}$ as the union of its left and right ones at level $\ell^{\prime}$, i.e., $\Gamma_{\ell^{\prime}}{(u)}=\Gamma_{\ell^{\prime}}^{L}(u)\cup\Gamma_{\ell^{\prime}}^{R}(u)$ and to its overall set of children as the union of each $\Gamma_{\ell^{\prime}}{(u)}$ for all $\ell^{\prime}<\ell$ , i.e., $\Gamma{(u)}=\bigcup_{\ell^{\prime}=0}^{\ell-1}\Gamma_{\ell^{\prime}}{(u)}$. Furthermore, we notice that two neighboring nodes $u$ and $w$ at level $\ell$ share a subset of their children. In a similar way, we introduce the concept of parent of a node in a skip list. Given a node $u$ at some level $\ell<\ell_{\max}(\mathcal{B})$, we say that $z$ is a parent of $u$ if $u\in\Gamma{(z)}$. Observe that each node $u$ such that $\ell_{\max}(u)<\ell_{\max}(\mathcal{B})$ has exactly two parents. Figure 8 shows an example of the children-parent relationship defined above. The children set of a node $u$ can be seen as a depth $1$ tree rooted in $u$ where $u$’s children are its leaves (see Figure 8(b)).
@@
The merge protocol is divided in two phases: (i) a preprocessing phase on the buffer network $\mathcal{B}$ and (ii) a merge phase on the clean network $\mathcal{C}$. The preprocessing phase takes care of creating the cohesive groups at each level that must be merged into $\mathcal{C}$ and to “shorten” the distance between every couple of nodes in such groups to guarantee fast communication time [^9] After the preprocessing phase, every couple of nodes in the same cohesive group can exchange messages within $\mathcal{O}(1)$ rounds. Thus, a node $u\in\mathcal{B}$ can send messages to its children $\Gamma(u)$ in $\mathcal{O}(1)$ time as well. within the cohesive groups. More precisely, let $X_{\ell}=\{x_{0},x_{1},x_{2},\dots,x_{k}\}$ be a set of consecutive nodes such that $\ell_{\max}^{\mathcal{B}}(x)=\ell$ for each $x\in X_{\ell}$ and $0\leq\ell\leq\ell_{\max}(\mathcal{B})$. Then, each node $x_{i}\in X_{\ell}$ for $1\leq i\leq k$ sends its ID to $x_{0}$ through its left neighbor. Doing so will allow each $x_{i}$ to discover the ID of all the nodes $x_{j}\in X_{\ell}$ such that $x_{j}>x_{i}$ (see Figure 9). We observe that the preprocessing phase requires $\mathcal{O}(\log n)$ rounds w.h.p.

### v2

@@ 2.4 Buffer Creation
To construct the Buffer Network $\mathcal{B}$, we begin with a set of nodes that recently joined the system due to churn. These nodes are initially unsorted and must be organized into a skip list structure. Here, the core challenge is to sort them quickly and reliably in a fully distributed setting. The current best-known approach to sort $n$ elements using $\text{polylog}{(n)}$ incoming and outgoing messages for each node during each round requires $\mathcal{O}(\log^{3}n)$ time [6]. To improve on such result, we rely on sorting networks theory (e.g. [16, 1, 53]). A sorting network is a fixed graph-like structure designed specifically for sorting any input permutation efficiently. More precisely, it can be viewed as a circuit-looking directed acyclic graph with $n$ inputs and $n$ outputs in which each node has exactly two inputs and two outputs (i.e., two incoming edges and two outgoing edges). The key property of a sorting network is that, for any input permutation, the outputs are guaranteed to be sorted. A sorting network is characterized by two parameters: the number of nodes resp. size and the depth. The space required to store a sorting network is determined by its size while the overall running time of the sorting algorithm is determined by its depth. Figure 7 shows an example of sorting network on $4$ elements.
@@
Figure 7: Example of a sorting network [16] on 44 numbers. Every crossing edge connecting two nodes in different levels is a node in the DAG representation. The upper output of a crossing edge connecting two nodes in different levels is the minimum between the two inputs, while the lower output is the maximum. Execution of the sorting network is left to right.
Here we use the AKS network [1] that allows to sort $n$ elements in $\mathcal{O}(\log n)$ rounds.
To use sorting networks in our distributed setting, we must simulate their structure on a suitable topology. In our case, this is achieved by embedding the AKS network onto a twin-butterfly, which we construct on top of the Spartan network.
Indeed, it is well known that sorting networks can be simulated on structured topologies such as a butterfly networks or other hypercubic networks (see Chapter 3.5 in [53]). To leverage this in our setting, we aim to embed the AKS sorting network onto a butterfly-like network. For this purpose, we use a result by Maggs and Vöcking [56] which shows that an AKS network can be embedded into a multibutterfly [83], a generalization of the classic butterfly topology.
A $k$-dimensional multibutterfly network consists of $k+1$ levels, each consisting of $2^{k}$ nodes. Nodes at level $\ell$ are labeled as $(\ell,i)$ for $0\leq\ell\leq k$ and $0\leq i<2^{k}$. Each level is partitioned into $2^{k}$ groups
$A_{\ell,j}=\left\{(\ell,i):\left\lfloor\frac{i}{2^{k-\ell}}\right\rfloor=j\right\}\quad\text{ for }0\leq j<2^{k}$
Nodes in group $A_{\ell,j}$ are connected to those in $A_{\ell+1,2j}$ and $A_{\ell+1,2j+1}$, creating a butterfly-like network.
To embed the AKS network, we focus on a specific subclass of multibutterfly networks called k-folded butterflies. These are obtained by superimposing $k$ butterfly networks and “folding” them into one. That is, the nodes from each butterfly with the same label are merged together, and the edges are colored so that each color class forms a valid butterfly. This folding reduces the number of physical nodes while preserving the logical connectivity of multiple butterflies.
Formally, suppose the edges of a $d$-dimensional multibutterfly of degree $4k$ can be colored by $k$ colors such that the network induced by the edges of each color are isomorphic to the $d$-dimensional butterfly. Then this multibutterfly can be constructed by folding $k$ butterfly networks. By folding, we mean that the labels of the nodes in the $A_{\ell,j}$ sets of each of these butterflies are permuted and the $k$ nodes with the same label in distinct butterflies are merged together to form a multibutterfly node [^7] Intuitively, a $k$-folded butterfly overlays $k$ butterfly networks and merges nodes with the same label, allowing multiple logical networks to coexist on a single physical graph..
In particular, Maggs and Vöcking [56] show how to embed the AKS network into a twin-butterfly, which is the case where two butterfly networks are folded together (i.e., a $2$-folded butterfly). In such a network, each leaf node has degree $4$, and each internal node has degree $8$. For completeness we state the result we are interested in.
@@
Here the load of an embedding is the maximum number of nodes of the AKS network mapped to any node of the twinbutterfly. The congestion is the maximum number of paths that use any edge in the twinbutterfly, and the dilation is the length of the longest path in the embedding.
@@
The idea is to use such a butterfly construction algorithm to build a twinbutterfly $\mathcal{M}$ in $\mathcal{O}(\log n)$ rounds and simulate the AKS algorithm on it. Moreover, $\mathcal{M}$ being a data structure build on top of the Spartan network, it naturally enjoys a churn resiliency property. In fact, every time a node $v$ is deleted from $\mathcal{S}$, a committee will temporary cover for $v$ in $\mathcal{M}$ (see Section 2.2). Finally, we show that the $\mathcal{B}$ can be created in $\mathcal{O}(\log n)$ rounds.
@@
$\mathcal{M}$ can be built in $\mathcal{O}(\log n)$ rounds (Lemma 2.6). Moreover, to handle cases in which $\kappa\cdot N+o(N)$ (see Theorem 2.5) is greater than the number of nodes in the Spartan network $\mathcal{S}$, we allow for nodes to represent a constant number of temporary “dummy” nodes for the sake of building the needed multibutterfly. Once $\mathcal{M}$ has been constructed, we run AKS on such network and after $\mathcal{O}(\log n_{\mathcal{B}})=\mathcal{O}(\log n)$ rounds we obtain the sorted list at the base of the new skip list we want to build. Next, each node $u$ in the buffer computes its maximum height $\ell_{\max}^{\mathcal{B}}(u)$ in the buffer skip list $\mathcal{B}$ and each level $0<\ell\leq\ell_{\max}(\mathcal{B})$ is created by copying the base level of the skip list. Moreover, each node $u$ at level $0<\ell\leq\ell_{\max}(\mathcal{B})$ can be of two types, effective or fill-in. We say that a node $u$ is effective at level $\ell$ if $\ell\leq\ell_{\max}^{\mathcal{B}}(u)$, and fill-in otherwise (see Figure 8(a)). Next, we use the same parallel $\mathcal{O}(\log n)$ rounds rewiring technique used in the Deletion phase (see Section 2.3) to exclude fill-in nodes from each level and obtain a skip list of effective nodes (see Figures 8(b)-8(c)). Wrapping up, to create the buffer network we need $\mathcal{O}(\log n)$ rounds to build the twibutterfly $\mathcal{M}$, $\mathcal{O}(\log n)$ rounds to run AKS on $\mathcal{M}$, $\mathcal{O}(\log n)$ rounds w.h.p. to duplicate the levels (see Lemma B.1 in Appendix B) and $\mathcal{O}(\log n)$ rounds to run the rewiring procedure. Thus, we need $\mathcal{O}(\log n)$ roundsw.h.p.
@@
This completes the creation of the Buffer network $\mathcal{B}$, which is now a properly structured skip list ready to be merged into the clean network $\mathcal{C}$.
@@ 2.5 Merge
Once the Buffer Network $\mathcal{B}$ is constructed, the next step is to merge it into the Clean Network $\mathcal{C}$, updating the distributed skip list to include the newly joined nodes. This operation must be carried out in a fully distributed manner and under high churn.
To this end, we introduce the WAVE protocol, a distributed merge algorithm based on a top-down traversal of $\mathcal{C}$ by groups of nodes in $\mathcal{B}$. The merge proceeds in rounds, with buffer nodes moving in cohesive groups, that are set of adjacent nodes at the same level in $\mathcal{B}$ that move and act together. These groups progressively insert themselves into the skip list $\mathcal{C}$, level by level, forming a “wave” of parallel insertions that sweeps from the top level down to the base level.
The merge phase consists of two main ingredients: (i) A preprocessing phase, executed on $\mathcal{B}$, and (ii) the WAVE merge protocol, executed on $\mathcal{C}$, in which the groups traverse and progressively integrate themselves into the skip list in parallel.
Cohesive Groups
A cohesive group $X=\{x_{0},x_{1},\dots,x_{k}\}\subseteq\mathcal{B}$ at some level $\ell$ is a set of contiguous nodes such that: (i) each node in $X$ has maximum height $\ell$, i.e., $\ell_{\max}^{\mathcal{B}}(x_{i})=\ell$, (ii) the group is sorted in increasing order.
Each group elects as a leader the node with the smallest ID (i.e., value) in the group, which will coordinate the merge actions on behalf of the group. Led by the leader, the cohesive group traverses $\mathcal{C}$ just like a node being inserted into a skip list. As the group moves through $\mathcal{C}$, it may split into subgroups because the cohesive group encountered a node $w$ in $\mathcal{C}$ with some elements of $X$ being smaller than $w$, while others are greater. Note that a split requires two rounds, one for the leader to inform all nodes in the new cohesive group who their leader is, and one for the nodes in the new cohesive group to “introduce” themselves to their new leader. We now observe a key structural property of how cohesive groups behave during traversal.
Proposition 2.10 (Cohesive Group Split Behavior).
Let $X=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{i-1}\leq x_{i}$ for $i\in[1,k]$ be a set of nodes that are traversing the skip list $\mathcal{C}$ all together (i.e., as a cohesive group). And let $v\in X$ be a node that separates from $X$ at some time $t\geq 0$, then all $u\in X$ such that $u\geq v$ will separate from $X$ as well, and will go along with $v$.
@@
For the sake of contradiction, let $X=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{i-1}\leq x_{i}$ for $i\in[1,k]$ be a cohesive group that is traversing the skip list. Moreover, assume that $x_{j}$ separates from $X$ at some time $t$, and that there exists $x_{\ell}$ for $\ell\geq j$ that does not follow $x_{j}$, i.e., $x_{\ell}$ does not separate from $X$. This generates a contradiction on the assumption that the elements in $X$ are sorted in ascending order.
Proposition 2.10 implies that when a cohesive group splits during its traversal of $\mathcal{C}$, the split always occurs at a single position along the ordered group. As a result, the group divides into exactly two contiguous subgroups: a left sub-group that continues the current traversal, and a right sub-group that separates and begins its own traversal (under a new leader). Therefore, at any round, a cohesive group can split into at most two new cohesive groups. Hence, we have that the traversal behavior of a cohesive group is tightly structured: in each round, the group either moves forward together, or splits cleanly into two contiguous subgroups. This ensures that group behavior remains predictable and parallelizable, with only one new leader elected per split. In particular, since group members perform the same operations up to the point of divergence, the cost of inserting a group is not significantly worse than inserting a single element. We formalize this next.
Lemma 2.12 (Cohesive Group Insertion Time Bound).
Let $X=\{x_{0},x_{1},\dots,x_{k}\}$ where $x_{i-1}\leq x_{i}$ for $i\in[1,k]$ be a set of cohesive nodes that are inserted all together at the same time and from the top left sentinel node in a skip list $\mathcal{C}$. Then, the time to insert the entire group $X$ is at most twice the time needed to insert just $x_{k}$.
@@
We prove the lemma by induction on the number of nodes in the set $X$. As a base case we consider $X=\{x_{0}\}$. Since is the only node, its insertion follows its optimal search path from the top-left sentinel in $\mathcal{C}$. Therefore, the insertion time for $x_{0}$ is exactly the same wether inserted alone or as a part of a singleton group. For the inductive hypothesis, we assume that for some $k\geq 1$, the time to insert the group $X=\{x_{0},x_{1},\dots,x_{k-1}\}$ is at most twice the time needed to insert only $x_{k-1}$, and that each element in the group follows its optimal search path when inserted from the top-left sentinel. For the inductive step we consider the insertion of $x_{k}$, along with the existing group $X=\{x_{0},x_{1},\dots,x_{k-1}\}$ (i.e., $X\cup\{x_{k}\}$). The key observation is that the search path of $x_{k}$ can be decomposed in two parts: (i) a shared prefix with $x_{k-1}$’s search path, and (ii) a unique suffix where $x_{k}$ diverges and continues alone towards its insertion point. Since the nodes move in parallel (cohesively) and are inserted from the same starting point, they make the same local decisions up to their divergence. By the inductive hypothesis, $x_{k-1}$’s search path is optimal, so the shared prefix must also be part of the optimal path for $x_{k}$. The remaining part (i.e., suffix) is also optimal by correctness of the skip list insertion algorithm [67]. Hence, the entire search path for $x_{k}$ is optimal. Since all insertions proceed in parallel, the total insertion time is determined by the longest search path among the group. By the inductive hypothesis, the insertion of $\{x_{0},x_{1},\dots,x_{k-1}\}$ takes at most two times the time to insert $x_{k-1}$ alone. Now we add $x_{k}>x_{k-1}$, whose path dominates $x_{k-1}$’s due to the group’s ordering. Since both, shared and unique, parts of $x_{k}$’s path are optimal, the total time remains within twice the time to insert $x_{k}$ alone.
This lemma ensures that even in the worst case, inserting a full group in parallel does not significantly increase the insertion time compared to a single node. Next, we examine how they are positioned within the skip list, and how the skip list’s structure ensures a clean separation between the levels. The following proposition captures two important structural properties of cohesive groups.
@@
Let $\ell$ be the maximum height of the skip list $\mathcal{B}$ and assume there are two disjoint cohesive groups $X_{\ell}$ and $Y_{\ell}$. This implies that $X_{\ell}$ and $Y_{\ell}$ are separated by one (or more) nodes $v$ such that $\ell_{\max}^{\mathcal{B}}(v)>\ell$. Thus we have a contradiction on $\ell$ being $\mathcal{B}$’s maximum height. Next, let $0\leq\ell<\ell_{\max}(\mathcal{B})$, consider the cohesive group $X_{\ell}=\{x_{0},x_{1},\dots,x_{k}\}$ and its left neighbor $v$ at level $\ell$ (i.e., $v\in N_{\mathcal{B}}^{(\texttt{left},\ell)}(x_{0})$). Then $v$ must be such that $\ell_{\max}^{\mathcal{B}}(v)>\ell$. That is because, if $\ell_{\max}^{\mathcal{B}}(v)=\ell$ then we would have $v\in X_{\ell}$ and if $\ell_{\max}^{\mathcal{B}}(v)<\ell$, we would have $v\notin N_{\mathcal{B}}^{(\texttt{left},\ell)}(x_{0})$.
These properties are crucial, they imply that every cohesive group is “encapsulated” between nodes of higher height, and in particular, that each group has two natural parents at the levels above.
Parent-Children Relationship in $\mathcal{B}$
To coordinate the merge across levels, we define a parent-child relationship among nodes in $\mathcal{B}$. Given a node $u$ at level $\ell>0$, its children at level $\ell^{\prime}<\ell$ are defined as the maximal contiguous block of nodes adjacent to $u$ at level $\ell^{\prime}$, whose heights are exactly $\ell^{\prime}$, and such that all nodes between them (if any) have height at most $\ell^{\prime}$. Formally, we define
$\Gamma_{\ell^{\prime}}^{\texttt{left}}(u)=\{v:v<u\land\ell_{\max}(v)=\ell^{\prime}\land\forall v<w<u,\ell_{\max}(w)\leq\ell^{\prime}\}$
and
$\Gamma_{\ell^{\prime}}^{\texttt{right}}(u)=\{v:v>u\land\ell_{\max}(v)=\ell^{\prime}\land\forall u<w<v,\ell_{\max}(w)\leq\ell^{\prime}\}$
as the left and right children of $u$ at level $\ell^{\prime}<\ell$ in $\mathcal{B}$, respectively. Moreover, we refer to $u$’s set of children at level $\ell^{\prime}$ as the union of its left and right ones at level $\ell^{\prime}$, i.e., $\Gamma_{\ell^{\prime}}{(u)}=\Gamma_{\ell^{\prime}}^{\texttt{left}}(u)\cup\Gamma_{\ell^{\prime}}^{\texttt{right}}(u)$ and to its overall set of children as the union of each $\Gamma_{\ell^{\prime}}{(u)}$ for all $\ell^{\prime}<\ell$ , i.e., $\Gamma{(u)}=\bigcup_{\ell^{\prime}=0}^{\ell-1}\Gamma_{\ell^{\prime}}{(u)}$. Furthermore, we notice that two neighboring nodes $u$ and $w$ at level $\ell_{\max}^{\mathcal{B}}$ share a subset of their children. Each node $u$ at level $\ell<\ell_{\max}^{\mathcal{B}}$ has two parents at level $\ell^{\prime}>\ell$ one to the left and one to the right. These parent nodes are responsible for notifying $u$ when it can safely initiate its own merge into $\mathcal{C}$ and when to update its pointers.
This relationship forms the backbone of the merge dependency graph, ensuring that merge operations flow top-down through the skip list levels in a pipelined [^8] By pipelined, we mean that they are executed one after the other. manner. It also allows nodes that are not in the merge phase yet to track the progress of their parents and initiate their own actions only when dependencies are resolved.
Figure 9(b) shows an example of the parent-children relationship in a skip list.
With this parent-children structure in place, the WAVE protocol guarantees that cohesive groups at lower levels only activate once their parents have successfully merged. This enables a clean pipelined traversal of $\mathcal{C}$ and maintains correctness and synchronization.
@@
The Preprocessing Phase in $\mathcal{B}$
Before the WAVE protocol begins, we perform a preprocessing phase in $\mathcal{B}$ to prepare each cohesive group for coordinated insertion. The phase includes:
Group Identification.
At each level $\ell$, we identify all maximal cohesive groups. This can be done locally by the nodes as follows: let $X_{\ell}=\{x_{0},x_{1},x_{2},\dots,x_{k}\}$ be a set of consecutive nodes such that $\ell_{\max}^{\mathcal{B}}(x)=\ell$ for each $x\in X_{\ell}$ and $0\leq\ell\leq\ell_{\max}(\mathcal{B})$. Then, each node $x_{i}\in X_{\ell}$ for $1\leq i\leq k$ sends its ID to $x_{0}$ through its left neighbor. Doing so will allow each $x_{i}$ to discover the ID of all the nodes $x_{j}\in X_{\ell}$ such that $x_{j}>x_{i}$ (see Figure 10).
Leader Election.
Within each group $X_{\ell}=\{x_{0},\dots,x_{k}\}$, the leftmost node $x_{0}$ elects itself as the group leader, and broadcasts this information to the other group members. This operation takes $\mathcal{O}(1)$ rounds since distances among members in the cohesive groups have been shortened in the previous step.
Parent Discovery.
Each node identifies its two parents (left and right) at the level(s) above by examining their pointers in $\mathcal{B}$. These parent links define the merge dependency graph used in the WAVE protocol.
State Initialization.
All nodes initialize their local state as idle except the leader on the single cohesive group in $\mathcal{B}$’s top-level, which are immediately activated to begin the merge.
Each step above can be performed independently at every level of $\mathcal{B}$, and each node only participates in one group. Since the number communication is local within groups, the preprocessing phase needs $\mathcal{O}(\log n)$ rounds with high probability in the first step. This follows from the fact that each cohesive group at some level $\ell$ contains at most $\mathcal{O}(\log n)$ nodes w.h.p. (see Lemma B.3 in Appendix B). After this step each node within the same group will be at distance one from all the other nodes in the group. This implies that the remaining steps can be performed in constant number of rounds. This is a key property since it allows nodes to communicate with their children and the other nodes within their same group in $\mathcal{O}(1)$ rounds.